Every story and comment from the Hacker News API, in a layout borrowed from 2001.
A comment on What is the fastest way for a programmer to make $10000?. Parent comment
Given the chance of winning a round is p=0.5, the chance of losing is q = 1 - p = 0.5
If you've lost the first round, the odds of you twice in a row are q = 0.5^(2 rounds) = 0.25
Which means the odds of you winning are p = 1 - q = 1 - 0.25 = 0.75.
Given the chance of winning a round is p=0.5, the chance of losing is q = 1 - p = 0.5
If you've lost the first round, the odds of you twice in a row are q = 0.5^(2 rounds) = 0.25
Which means the odds of you winning are p = 1 - q = 1 - 0.25 = 0.75.